OpenStudy (anonymous):

sec^2 (pi/5) - tan^2 (pi/5)

OpenStudy (jdoe0001):

noted... thanks for sharing that

myininaya (myininaya):

oh ok do you know your Pythagorean identities?

OpenStudy (anonymous):

sin ^2 + cos ^2 = 1 i know theres 3 more but i forgot :/

myininaya (myininaya):

you can get the others by dividing that equation by cos^2 or dividing it by sin^2 let's just divide both sides by cos^2 for now \[\sin^2(x)+\cos^2(x)=1 \\ \frac{\sin^2(x)+\cos^2(x)}{\cos^2(x)}=\frac{1}{\cos^2(x)} \\ \tan^2(x)+1=\sec^2(x) \\ 1=\sec^2(x)-\tan^2(x)\]

OpenStudy (anonymous):

so my problem = 1?

myininaya (myininaya):

yep

OpenStudy (anonymous):

wow i have 1 more im also confused on

myininaya (myininaya):

k what is it

OpenStudy (anonymous):

\[(\cos 2\pi/9)(\sec 2\pi/9)\]

myininaya (myininaya):

cosine and secant are reciprocals of one another

myininaya (myininaya):

\[\sec(x)=\frac{1}{\cos(x)}\]

OpenStudy (anonymous):

oh yeah! so it = 1 lol

myininaya (myininaya):

yes

OpenStudy (anonymous):

im so done....plus i never got taught those identities yet

myininaya (myininaya):

what is this? like itt?

myininaya (myininaya):

like college?

OpenStudy (anonymous):

im in high school precalc honors

myininaya (myininaya):

ah

myininaya (myininaya):

well i think in this is my opinion teachers shouldn't have to teach everything

myininaya (myininaya):

just the basics and then students can use those basics to make their own discoveries

OpenStudy (anonymous):

I also couldn't find the examples in the book

OpenStudy (anonymous):

to help solve for those 2

myininaya (myininaya):

well now you have them examples that is

OpenStudy (anonymous):

thank you